# Convergence Rates for the Maximum A Posteriori Estimator in PDE-Regression Models with Random Design

### v1

@@ Appendix A Appendix
We now approaching the proof of Theorem 2.6. Therefore, we need some notation and further helping lemma.
Notation § A.5 (Empirical Measures).
For $N\in\mathbb{N}$ let $D_{N}:=(Y_{i},Z_{i})_{i\in\left\llbracket N\right\rrbracket}\in(V\times\mathcal{Z})^{N}$ be a data set. For any $(y,z)\in V\times\mathcal{Z}$ denote by $\delta_{(y,z)}(\cdot)$ the Dirac measure on $V\times\mathcal{Z}$. The empirical measure on $(V\times\mathcal{Z},\mathscr{B}_{V\times\mathcal{Z}})$ based on $D_{N}$ is then defined by
^$\widehat{\mathbb{P}}_{N}(A):=\frac{1}{N}\sum_{i\in\left\llbracket N\right\rrbracket}\delta_{(Y_{i},Z_{i})}(A)$
for any $A\in\mathscr{B}_{V\times\mathcal{Z}}$. Evidently, we have for any measurable function $h:V\times\mathcal{Z}\to\mathbb{R}$
@@
The following helping lemma provides M-estimation techniques, which are used to prove Theorem 2.6. The proof follows along the lines of van de Geer [2001], Nickl et al. [2020] and Bohr and Nickl [2023].
Lemma § A.6 (M-estimation techniques).
With the notation of Theorem 2.6 and the lines above it, we have for any $\theta_{\mathfrak{o}}\in\tilde{\Theta}$ and $\tilde{\theta}\in\Theta$, $\mathfrak{r}\in\mathbb{R}_{>0}$ and any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$
$\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\leq 2\int_{V\times\mathcal{Z}}\log\left(\frac{p_{\hat{\theta}_{\textsc{MAP}}}(y,z)}{p_{\tilde{\theta}}(y,z)}\right)\mathrm{d}(\widehat{\mathbb{P}}_{N}-\mathbb{P}_{\theta_{\mathfrak{o}}})(y,z)+\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})$ (A.2)
@@
On $\Xi_{\mathfrak{r},R}$, we have further
$\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta})\geq c_{1}\cdot R^{2}$ (A.3)
@@
$\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})-\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})\geq\frac{1}{6}\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta}).$ (A.4)
Proof of 5 (Lemma A.6).
The proof of Eq. A.2 is elementary and follows along the lines of Lemma 4.10 in Nickl and Wang [2024] (replacing the finite-dimensional approximation of $\theta_{\mathfrak{o}}$ by some arbitrary $\tilde{\theta}\in\Theta$). We continue with the proof of Eq. A.3. On the event $\Xi_{\mathfrak{r},R}$, we have
@@
$\leq 2\left(\frac{1}{2}\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\ -\left\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\right)+\mathfrak{r}^{2}\left(\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}-2\left\|\tilde{\theta}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}\right)$
$\leq 2\left(\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}-\left\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\right)^{2}+2\mathfrak{r}^{2}\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}-\mathfrak{r}^{2}\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}-2\mathfrak{r}^{2}\left\|\tilde{\theta}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}$
$\leq 2\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta}),$
where we have used $\frac{1}{2}a^{2}-b^{2}\leq(a-b)^{2}$ for two real numbers $a,b\in\mathbb{R}$ as well as the reversed triangle inequality. Rearranging, the last inequality shows Eq. A.3. Eq. A.4 follows by the same elementary computations as in van de Geer [2001] (see pp. 3-4).
Proof of 6 (Theorem 2.6).
The proof follows the key ideas of Bohr and Nickl [2023]. Since we are allowing a more general $\tilde{\theta}\in\Theta$, we also need to adapt the ideas of Nickl et al. [2020]. We start with the exponential inequality Eq. 2.4. Fix $\tilde{\theta}\in\Theta$. According to Lemma A.6, it sufficies to controll the $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$-probability of the event
@@
and obtain with Eq. A.3 and Eq. A.4 from Lemma A.6
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}(\tilde{\Xi}_{\mathfrak{r},R})\leq\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta})\geq c_{1}^{2}R^{2},\,\frac{1}{12}\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta})\leq|\mathcal{W}_{N}(\hat{\theta}_{\textsc{MAP}})|\right\}\right).$
for some $c_{1}\in\mathbb{R}_{\geq 1}$. In the following step, we apply a so-called peeling device, where we refer to van de Geer [2000] for a generic explanatation. This method was already used in a similar way in Nickl et al. [2020] and Bohr and Nickl [2023] to controll the process $\mathcal{W}_{N}(\hat{\theta}_{\textsc{MAP}})$ over disjoint slices of a subset of $\Theta$. In detail, for all $l\in\mathbb{N}$ and $c_{1}$ sufficiently large, define the family of slices
@@
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\Xi_{\mathfrak{r},R}\right)\leq\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}|\mathcal{W}_{N}(\theta)|\geq\frac{2^{2l}}{48}R^{2}c_{1}^{2}\right\}\right).$
We follow the approach of Bohr and Nickl [2023] and decompose $\mathcal{W}_{N}$ into two empirical processes. Indeed, for each $\theta\in\Theta$, simple computations show
$N\mathcal{W}_{N}(\theta)$ $=\ell_{N}(\theta)-\ell_{N}(\tilde{\theta})-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[\ell_{N}(\theta)-\ell_{N}(\tilde{\theta})],$
which evidently defines a centered process. Given the data $D_{N}\sim\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$, we further decompose the log-likelihood difference of the last display by
$\ell_{N}(\theta)-\ell_{N}(\tilde{\theta})$ $=\frac{1}{2\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left\{\left|Y_{i}-\mathscr{G}(\tilde{\theta})(Z_{i})\right|_{V}^{2}-\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}\right\}$
@@
$T_{1}(\theta):=\frac{1}{\sigma\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\langle\varepsilon_{i},\mathscr{G}(\theta)(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})\rangle_{V},$ (A.5)
@@
$T_{2}(\theta):=\frac{1}{2\sigma^{2}\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\left\{\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}\right\},$ (A.6)
respectively, we have that $\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[T_{1}(\theta)]=0$, as $Z$ and $\varepsilon$ are independent and $\varepsilon$ is centered. In other words, we have derived so far
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\Xi_{\mathfrak{r},R}\right)$ $\leq\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}|\mathcal{W}_{N}(\theta)|\geq\frac{2^{2l}}{48}R^{2}c_{1}^{2}\right\}\right)$
$\leq\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}\left\{|T_{1}(\theta)|\right\}\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right\}\right)+\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}\left\{|T_{2}(\theta)-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[T_{2}(\theta)]|\right\}\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right\}\right).$
We will discuss the two different probabilties separately and starting with the the multiplier-type process defined in Eq. A.5. To apply Lemma A.7, we denote for each $j\in\left\llbracket d_{V}\right\rrbracket$ by $\mathscr{G}(\cdot)_{j}:\Theta\to\mathbb{L}^{2}(\mathcal{O})$ the $j$-th coordinate of the forward map $\mathscr{G}:\Theta\to\mathbb{L}^{2}(\mathcal{O},V)$, where we again identify $V$ with $\mathbb{R}^{d_{V}}$. We define a class of measurable functions by
$\mathcal{H}_{l,j}:=\left\{\mathscr{G}(\tilde{\theta})_{j}-\mathscr{G}(\theta)_{j}:\mathcal{Z}\to\mathbb{R}:\,\theta\in\Theta_{l}\right\}.$ (A.7)
To shorten notation in the subsequent, we write $h_{\theta,j}:=\mathscr{G}(\theta_{\mathfrak{o}})_{j}-\mathscr{G}(\theta)_{j}$. Thus, for any $\theta\in\Theta$, we have
[…]
where $\varepsilon_{ij}\overset{i.i.d.}{\sim}\operatorname{N}(0,1)$ for $(i,j)\in\left\llbracket N\right\rrbracket\times\left\llbracket d_{V}\right\rrbracket$. Conclusevly, we see
@@
We now verify that the function class $\mathcal{H}_{l,j}$ defined in Eq. A.7 satisfies the assumptions formulated in Lemma A.7. Indeed, by C2) in 2.3, we have uniformly for all $\theta\in\Theta_{l}$
$\sup_{z\in\mathcal{Z}}\left\{|h_{\theta,j}(z)|\right\}=\sup_{z\in\mathcal{Z}}\left\{|\mathscr{G}(\tilde{\theta})_{j}(z)-\mathscr{G}(\theta)_{j}(z)|\right\}\leq\sup_{z\in\mathcal{Z}}\left\{\left|\mathscr{G}(\tilde{\theta})(z)-\mathscr{G}(\theta)(z)\right|_{V}\right\}\leq 2C_{\mathscr{G},\operatorname{B}}.$
@@
$\mathbb{E}^{Z}[h_{\theta,j}^{2}(Z)]\leq\sum_{j\in\left\llbracket d_{V}\right\rrbracket}\int_{\mathcal{Z}}\left|\mathscr{G}(\theta)_{j}(Z)-\mathscr{G}(\tilde{\theta})_{j}(Z)\right|_{V}^{2}\mathrm{d}\zeta(z)=\left\|\mathscr{G}(\theta)-\mathscr{G}(\tilde{\theta})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}.$
@@
$\left\|\mathscr{G}(\theta)-\mathscr{G}(\tilde{\theta})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq\mathfrak{d}_{\mathfrak{r}}^{2}(\theta,\tilde{\theta})\leq 2^{2l}R^{2}c_{1}^{2}=:v_{l}^{2}.$
@@
where $d_{2}$ and $d_{\infty}$ are defined according to Lemma A.7. Starting with the first one, let $\theta_{1},\theta_{2}\in\Theta_{l}$ be arbitrarily chosen. Exploiting C1) in 2.3, we have
$d_{2}(\theta_{1},\theta_{2})=\left\|h_{\theta_{1},j}-h_{\theta_{2},j}\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}$ $\leq C_{\operatorname{Lip},2}\cdot\left(1+\left\|\theta_{1}\right\|^{\gamma_{1}}_{H^{\alpha}(\mathcal{O},W)}\lor\left\|\theta_{2}\right\|^{\gamma_{1}}_{H^{\alpha}(\mathcal{O},W)}\right)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{(H^{\kappa}(\mathcal{O},W))^{*}}.$
By definition of $\Theta_{l}$, $\theta_{1}$, $\theta_{2}$ are contained in a $H^{\alpha}$-ball, such that the polynomial growth in the last display can be upper bounded. It suffices to find a $\rho$-covering of $\Theta_{l}$ with respect to the dual $(H^{\kappa}(\mathcal{O},W))^{*}$-norm. Applying Lemma A.9, there exists a constant $c_{2}=c_{2}(\alpha,\kappa,d)\in\mathbb{R}_{>0}$, such that for all $\rho\in\mathbb{R}_{>0}$
$N(\Theta_{l},\left\|\cdot\right\|_{(H^{\kappa}(\mathcal{O},W))^{*}},\rho)\leq\exp\left(c_{2}\cdot d_{W}\left(\frac{2^{l}R\sqrt{d_{W}}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha+\kappa}}\right).$
Consequently, for every $\rho\in\mathbb{R}_{>0}$ there exists a $\rho$-covering $\{\theta^{(i)}\}_{i\in\left\llbracket N\right\rrbracket}$ of $\Theta_{l}$ with $N$ upper bounded by the line above. Thus, for all $\theta\in\Theta_{l}$, there exists some $i_{\circ}\in\left\llbracket N\right\rrbracket$ with
$d_{2}(\theta,\theta^{(i_{\circ})})=\left\|h_{\theta,j}-h_{\theta^{(i_{\circ})},j}\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}$ $\leq C_{\operatorname{Lip},2}\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{1}}\right)\cdot\rho=:m_{1}\cdot\rho.$
Thus,

### v3

@@ 4.1 Proofs of section 2
We are now approaching the proof of theorem 2.6. Therefore, we need some notations and further helping lemma.
Notation § 4.5 (Empirical Measures).
For $N\in\mathbb{N}$, let $D_{N}:=(Y_{i},Z_{i})_{i\in\left\llbracket N\right\rrbracket}\in(V\times\mathcal{Z})^{N}$ be given data. For any $(y,z)\in V\times\mathcal{Z}$, denote by $\delta_{(y,z)}(\cdot)$ the Dirac measure on $(V\times\mathcal{Z},\mathscr{B}_{V\times\mathcal{Z}})$ in $(y,z)$. The empirical measure on $(V\times\mathcal{Z},\mathscr{B}_{V\times\mathcal{Z}})$ based on $D_{N}$ is then for any $A\in\mathscr{B}_{V\times\mathcal{Z}}$ defined by
$\widehat{\mathbb{P}}_{N}(A):=\frac{1}{N}\sum_{i\in\left\llbracket N\right\rrbracket}\delta_{(Y_{i},Z_{i})}(A).$
Evidently, we have for any measurable function $h:V\times\mathcal{Z}\to\mathbb{R}$
@@
The following Lemma provides M-estimation techniques, which are used to prove theorem 2.6. The proof follows along the lines of van de Geer (2001), Nickl et al. (2020) and Bohr and Nickl (2023).
Lemma § 4.6 (M-Estimation Techniques).
With the notation of theorem 2.6 and the lines above it, for any $\theta_{\mathfrak{o}}\in\tilde{\Theta}$, $\tilde{\theta}\in\Theta$, $\mathfrak{r}\in\mathbb{R}_{>0}$ and any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$, we have
$\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\leq 2\int_{V\times\mathcal{Z}}\log\left(\frac{p_{\hat{\theta}_{\textsc{MAP}}}(y,z)}{p_{\tilde{\theta}}(y,z)}\right)\mathrm{d}(\widehat{\mathbb{P}}_{N}-\mathbb{P}_{\theta_{\mathfrak{o}}})(y,z)+\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})$ (18)
@@
On $\Xi_{\mathfrak{r},R}$, we further have
$\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta})\geq c_{1}\cdot R^{2}$ (19)
@@
[…]
Proof of lemma 4.6.
The proof of eq. 18 is elementary and follows along the lines of Lemma 4.10 in Nickl and Wang (2024) (replacing the finite-dimensional approximation of $\theta_{\mathfrak{o}}$ by some arbitrary $\tilde{\theta}\in\Theta$). We continue with the proof of eq. 19. On the event $\Xi_{\mathfrak{r},R}$, we have
@@
$\leq 2\left(\frac{1}{2}\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}-\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta_{\mathfrak{o}})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\right)$
$\hskip 170.71652pt+\mathfrak{r}^{2}\left(\|\hat{\theta}_{\textsc{MAP}}\|_{H^{\alpha}(\mathcal{O},W)}^{2}-2\|\tilde{\theta}\|_{H^{\alpha}(\mathcal{O},W)}^{2}\right)$
$\leq 2\left(\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}-\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta_{\mathfrak{o}})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\right)^{2}$
$\hskip 113.81102pt+2\mathfrak{r}^{2}\|\hat{\theta}_{\textsc{MAP}}\|_{H^{\alpha}(\mathcal{O},W)}^{2}-\mathfrak{r}^{2}\|\hat{\theta}_{\textsc{MAP}}\|_{H^{\alpha}(\mathcal{O},W)}^{2}-2\mathfrak{r}^{2}\|\tilde{\theta}\|_{H^{\alpha}(\mathcal{O},W)}^{2}$
[…]
where we have used $\frac{1}{2}a^{2}-b^{2}\leq(a-b)^{2}$ for two real numbers $a,b\in\mathbb{R}$ as well as the reversed triangle inequality. Rearranging, the last inequality shows eq. 19. eq. 20 follows by the same elementary computations as in van de Geer (2001) (see pp. 3-4). ∎
Proof of theorem 2.6.
The proof follows the key ideas of Bohr and Nickl (2023). Since we are allowing a general $\tilde{\theta}\in\Theta$, we also need to adapt the ideas of Nickl et al. (2020). We start with the exponential inequality eq. 11. Fix $\tilde{\theta}\in\Theta$. According to lemma 4.6, it sufficies to control the $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$-probability of the event
@@
and obtain with eq. 19 and eq. 20 from lemma 4.6
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}(\tilde{\Xi}_{\mathfrak{r},R})\leq\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(|\mathcal{W}_{N}(\hat{\theta}_{\textsc{MAP}})|\geq\frac{1}{12}\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta}),\,\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta})\geq c_{1}^{2}R^{2},\right).$
for some $c_{1}\in\mathbb{R}_{\geq 1}$. In the following step, we apply a so-called peeling device, where we refer to van de Geer (2000) for a generic explanation. This method was already used in a similar way in Nickl et al. (2020) and Bohr and Nickl (2023) to control the process $\mathcal{W}_{N}(\hat{\theta}_{\textsc{MAP}})$ over disjoint slices of $\Theta$. In detail, for all $l\in\mathbb{N}$ and $c_{1}$ sufficiently large, define the family of slices
$\Theta_{l}:=\left\{\theta\in\Theta:\,2^{2(l-1)}R^{2}c_{1}^{2}\leq\mathfrak{d}_{\mathfrak{r}}^{2}(\theta,\tilde{\theta})<2^{2l}R^{2}c_{1}^{2}\right\}\subseteq H^{\alpha}(\mathcal{O},W,2^{l}R\mathfrak{r}^{-1}c_{1}),$
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\Xi_{\mathfrak{r},R}\right)\leq\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|\mathcal{W}_{N}(\theta)|\geq\frac{2^{2l}}{48}R^{2}c_{1}^{2}\right).$
We follow the approach of Bohr and Nickl (2023) and decompose $\mathcal{W}_{N}$ into two empirical processes. Indeed, for each $\theta\in\Theta$, simple computations show
@@
which evidently defines a centered stochastic process. Given the data $D_{N}\sim\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$, we further decompose the log-likelihood difference of the last display by
@@
$\overset{d}{=}\frac{1}{\sigma}\sum_{i\in\left\llbracket N\right\rrbracket}\langle\varepsilon_{i},\mathscr{G}(\theta)(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})\rangle_{V}$
$\phantom{==}+\frac{1}{2\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left\{|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})|_{V}^{2}-|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\theta)(Z_{i})|_{V}^{2}\right\}.$
@@
$T_{1}(\theta):=\frac{1}{\sigma\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\langle\varepsilon_{i},\mathscr{G}(\theta)(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})\rangle_{V},$ (21)
@@
$T_{2}(\theta):=\frac{1}{2\sigma^{2}\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\left\{|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})|_{V}^{2}-|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\theta)(Z_{i})|_{V}^{2}\right\},$ (22)
respectively, we have that $\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[T_{1}(\theta)]=0$, as $Z$ and $\varepsilon$ are independent and $\varepsilon$ is centered. With these, we have derived so far
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\Xi_{\mathfrak{r},R}\right)$ $\leq\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|\mathcal{W}_{N}(\theta)|\geq\frac{2^{2l}}{48}R^{2}c_{1}^{2}\right)$
$\leq\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}\left\{|T_{1}(\theta)|\right\}\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right)$
$\phantom{==}+\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}\left\{|T_{2}(\theta)-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[T_{2}(\theta)]|\right\}\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right).$
We will discuss the two different probabilities separately and starting with the multiplier-type process defined in eq. 21. In order to apply lemma 4.7, we denote for each $j\in\left\llbracket d_{V}\right\rrbracket$ by $\mathscr{G}(\cdot)_{j}:\Theta\to\mathbb{L}^{2}(\mathcal{O})$ the $j$-th coordinate of the forward map $\mathscr{G}:\Theta\to\mathbb{L}^{2}(\mathcal{O},V)$, where we again identify $V$ with $\mathbb{R}^{d_{V}}$. We define a class of measurable functions by
$\mathcal{H}_{l,j}:=\left\{\mathscr{G}(\tilde{\theta})_{j}-\mathscr{G}(\theta)_{j}:\mathcal{Z}\to\mathbb{R}:\,\theta\in\Theta_{l}\right\}.$ (23)
To shorten notation in the subsequent, we write $h_{\theta,j}:=\mathscr{G}(\tilde{\theta})_{j}-\mathscr{G}(\theta)_{j}$. Thus, for any $\theta\in\Theta$, we have
$T_{1}(\theta)$ $\overset{d}{=}-\frac{1}{\sigma\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\sum_{j\in\left\llbracket d_{V}\right\rrbracket}(\mathscr{G}(\tilde{\theta})_{j}(Z_{i})-\mathscr{G}(\theta)_{j}(Z_{i}))\varepsilon_{ij}$
$=-\frac{1}{\sigma\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\sum_{j\in\left\llbracket d_{V}\right\rrbracket}h_{\theta,j}(Z_{i})\varepsilon_{ij}=:-\frac{1}{\sigma}\sum_{j\in\left\llbracket d_{V}\right\rrbracket}Z_{N,j}^{(1)}(\theta),$
where $\varepsilon_{ij}\overset{i.i.d.}{\sim}\operatorname{N}(0,1)$ for $(i,j)\in\left\llbracket N\right\rrbracket\times\left\llbracket d_{V}\right\rrbracket$. Conclusively, we see
@@
We now verify that the function class $\mathcal{H}_{l,j}$ defined in eq. 23 satisfies the assumptions formulated in lemma 4.7. Indeed, by C2) in 2.3, we have uniformly for all $\theta\in\Theta_{l}$
$\sup_{z\in\mathcal{Z}}\left\{|h_{\theta,j}(z)|\right\}=\sup_{z\in\mathcal{Z}}\left\{|\mathscr{G}(\tilde{\theta})_{j}(z)-\mathscr{G}(\theta)_{j}(z)|\right\}\leq\sup_{z\in\mathcal{Z}}\left\{|\mathscr{G}(\tilde{\theta})(z)-\mathscr{G}(\theta)(z)|_{V}\right\}\leq 2C_{\mathscr{G},\operatorname{B}}.$
@@
$\mathbb{E}^{Z}[h_{\theta,j}^{2}(Z)]\leq\sum_{j\in\left\llbracket d_{V}\right\rrbracket}\int_{\mathcal{Z}}|\mathscr{G}(\theta)_{j}(z)-\mathscr{G}(\tilde{\theta})_{j}(z)|_{V}^{2}\mathrm{d}\zeta(z)=\|\mathscr{G}(\theta)-\mathscr{G}(\tilde{\theta})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}.$
@@
$\|\mathscr{G}(\theta)-\mathscr{G}(\tilde{\theta})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq\mathfrak{d}_{\mathfrak{r}}^{2}(\theta,\tilde{\theta})\leq 2^{2l}R^{2}c_{1}^{2}=:v_{l}^{2}.$ (24)
@@
where $d_{2}$ and $d_{\infty}$ are defined according to lemma 4.7. Starting with the first one, let $\theta_{1},\theta_{2}\in\Theta_{l}$ be arbitrarily chosen. Exploiting C1) in 2.3, we have
$d_{2}(\theta_{1},\theta_{2})=\|h_{\theta_{1},j}-h_{\theta_{2},j}\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}\hskip-2.15277pt\leq C_{\operatorname{Lip},2}\left(1+\|\theta_{1}\|^{\gamma_{1}}_{H^{\alpha}(\mathcal{O},W)}\hskip-4.30554pt\lor\|\theta_{2}\|^{\gamma_{1}}_{H^{\alpha}(\mathcal{O},W)}\right)\hskip-2.15277pt\|\theta_{1}-\theta_{2}\|_{(H^{\kappa}(\mathcal{O},W))^{*}}.$
By definition of $\Theta_{l}$, $\theta_{1}$, $\theta_{2}$ are contained in a $H^{\alpha}$-ball, such that the polynomial growth in the last display can be upper bounded by the corresponding radius. Hence, it suffices to find a $\rho$-covering of $\Theta_{l}$ with respect to the dual $(H^{\kappa}(\mathcal{O},W))^{*}$-norm. Applying lemma 4.9 i), there exists a constant $c_{2}=c_{2}(\alpha,\kappa,d)\in\mathbb{R}_{>0}$, such that for all $\rho\in\mathbb{R}_{>0}$
$N(\Theta_{l},\|\cdot\|_{(H^{\kappa}(\mathcal{O},W))^{*}},\rho)\leq\exp\left(c_{2}\cdot d_{W}\left(\frac{2^{l}R\sqrt{d_{W}}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha+\kappa}}\right).$
As a conseqence, for every $\rho\in\mathbb{R}_{>0}$, there exists a $\rho$-covering $\{\theta^{(i)}\}_{i\in\left\llbracket N\right\rrbracket}$ of $\Theta_{l}$ with respect to $\|\cdot\|_{(H^{\kappa}(\mathcal{O},W))^{*}}$ with $N=N(\Theta_{l},\|\cdot\|_{(H^{\kappa}(\mathcal{O},W))^{*}},\rho)$. Hence, for all $\theta\in\Theta_{l}$, there exists some $i_{\circ}\in\left\llbracket N\right\rrbracket$ with
$d_{2}(\theta,\theta^{(i_{\circ})})=\|h_{\theta,j}-h_{\theta^{(i_{\circ})},j}\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}$ $\leq C_{\operatorname{Lip},2}\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{1}}\right)\cdot\rho=:m_{1}\cdot\rho$
and we deduce